# Learning Visual Quadrupedal Loco-Manipulation from Demonstrations

### v1

@@ II-C Loco-Manipulation
Manipulation skills for legged robots have been greatly improved recently. Some researchers choose to augment robots with additional hardware on their backs [27, 28, 29, 30], which significantly increases costs and the additional weight compromises the robot’s locomotion abilities. Others try to use the robot’s legs to perform manipulation tasks. However, these methods have several drawbacks and limitations. (i) The methods are designed for specific tasks, often restricted to predefined sequences of actions [1, 2, 3, 4]. In contrast, we propose a general framework in which a single agent is trained to solve a series of tasks; (ii) The manipulation is very coarse and does not allow for fine-grained manipulation [2, 4]. By contrast, our method can train a low-level controller to achieve precise control for the legs; (iii) Performing static manipulation, failing to leverage the inherent dynamic capabilities of legged robots [3]. In contrast, our method enables robots to carry out manipulation tasks while preserving their locomotion abilities; (iv) It is limited to 3-DoF point tracking, which falls short for more complex manipulation tasks. Additionally, the authors only implemented a low-level controller, necessitating human teleoperation for the execution of all tasks [31]. Conversely, our method, utilizing a visual planner, empowers the robot to autonomously execute complex daily manipulation tasks, by tracking both trajectories and orientations.
@@ III-A Overview
We develop a low-level controller $\pi_{c}$, which allows the robot to track a 6-DoF target trajectory with any forelimb while maintaining stable walking. This is achieved by training a policy whose inputs include the manipulator flag, the current desired end-effector pose, the current target point of the end-effector calculated by the rational Bézier curve, and the next three target points. The output is desired robot joint positions $q^{d}_{m}\in\mathbb{R}^{12}$. This control policy runs at $50$ Hz, and the desired torque $\mathbf{\tau}$ of the robot is obtained through a PD controller from the desired joint positions.
@@ III-C2 Expert Demonstration Collection
As illustrated in Fig. 3, we designed expert trajectory parameters ${}^{o}p$ for each task. To enable the robot to learn to manipulate objects placed in different poses, we randomized the positions and yaw axis angles of the objects. The expert trajectory parameters are represented in the object frame, allowing them to accurately follow the object. During the collection of expert demonstrations, the object frame is transformed to the world frame based on the object’s pose.
We gather expert data by having robots follow expertly designed trajectories, utilizing large-scale parallel simulations in IsaacGym [36]. This method enables the rapid collection of a significant volume of data. The collection time does not increase with the volume of data; it primarily depends on the size of the Video Random Access Memory (VRAM). In this paper, we collect 200 timesteps of expert demonstration data for each task, with 100 trajectories collected per task, taking approximately three minutes. The success rates of the data collected are shown in Tab. III, indicating that expert demonstrations collected by our method achieve relatively high success rates across most tasks.
Despite the collection of expert data in simulations, our method achieves seamless sim-to-real deployment in actual environments, thanks to the carefully designed hierarchical structure and the incorporation of point clouds as visual input.
@@ V-A Performance Comparison
Success rate (%) Expert HRL Ours
Press Button $93.67\scriptstyle{\pm 1.53}$ $21.67\scriptstyle{\pm 3.51}$ $\mathbf{92.33\scriptstyle{\pm 8.96}}$
Pull Handle $86.33\scriptstyle{\pm 3.21}$ $0.00\scriptstyle{\pm 0.00}$ $\mathbf{82.33\scriptstyle{\pm 3.21}}$
Push Door $88.00\scriptstyle{\pm 7.76}$ $15.67\scriptstyle{\pm 3.51}$ $\mathbf{85.33\scriptstyle{\pm 5.03}}$
Lift Basket $67.00\scriptstyle{\pm 9.54}$ $11.00\scriptstyle{\pm 4.36}$ $\mathbf{59.33\scriptstyle{\pm 12.01}}$
Open Dishwasher $20.67\scriptstyle{\pm 4.62}$ $1.33\scriptstyle{\pm 1.15}$ $\mathbf{5.67\scriptstyle{\pm 5.51}}$
Close Dishwasher $92.33\scriptstyle{\pm 5.51}$ $4.67\scriptstyle{\pm 1.53}$ $\mathbf{50.33\scriptstyle{\pm 8.14}}$
Pull Objects $42.00\scriptstyle{\pm 1.73}$ $8.00\scriptstyle{\pm 2.65}$ $\mathbf{12.67\scriptstyle{\pm 10.79}}$
Twist Valve $70.33\scriptstyle{\pm 2.31}$ $10.00\scriptstyle{\pm 4.58}$ $\mathbf{52.33\scriptstyle{\pm 10.21}}$
Shoot Football $37.67\scriptstyle{\pm 18.18}$ $3.67\scriptstyle{\pm 2.08}$ $\mathbf{26.00\scriptstyle{\pm 2.00}}$
@@ V-C Sim2real Transfer
We directly implement the trained controller and planner on a real-world Aliengo quadruped robot without the need for further fine-tuning, demonstrating the robustness of our approach. As shown in Fig. 2, the pose of the robot is determined using AprilTag for the coordinate system transformation of trajectory parameters and point clouds. Following this, we applied preprocessing to the point cloud to ensure alignment with the simulation. In Fig. 1, we illustrate the robot performing tasks like lifting a basket with its forelimbs and pushing the door during locomotion. For demonstrations of all the skills developed through our method, please refer to the accompanying video.
@@ VI Conclusion and future works
While we have demonstrated the effectiveness of our method in both simulation and real-world scenarios, there are areas for improvement. (i) The inference speed limitations of diffusion-based BC hinder task performance, making it challenging for robots to handle dynamic environments. (ii) While our approach to gathering expert demonstrations is significantly efficient, it does not robustly handle unexpected situations. Additionally, the diversity of the expert demonstrations collected is limited.
@@
This study represents a novel effort to master whole-body loco-manipulation, possessing boundless potential. (i) It showcases remarkable scalability, enabling rapid collection of expert data for either scale up or data-mixture with real-world data. (ii) To further bridge the simulation-to-reality gap, exploring the use of teleoperation for real-world data collection is an option. Given our method’s high data efficiency, this method could substantially lessen the required effort.

### v2

@@ II-C Loco-Manipulation
Manipulation skills for legged robots have been greatly improved recently. Some researchers choose to augment robots with additional hardware on their backs [27, 28, 29, 30], which significantly increases costs and the additional weight compromises the robot’s locomotion abilities. Others try to use the robot’s legs to perform manipulation tasks. However, these methods have several drawbacks and limitations. (i) The methods are designed for specific tasks, often restricted to predefined sequences of actions [1, 2, 3, 4]. In contrast, we propose a general framework in which a single agent is trained to solve a series of tasks; (ii) The manipulation is very coarse and does not allow for fine-grained manipulation [2, 4]. By contrast, our method can train a low-level controller to achieve precise control for the legs; (iii) Performing static manipulation, failing to leverage the inherent dynamic capabilities of legged robots [3]. In contrast, our method enables robots to carry out manipulation tasks while preserving their locomotion abilities; (iv) It is limited to 3-DoF point tracking, which falls short for more complex manipulation tasks. Additionally, the authors only implemented a low-level controller, necessitating human teleoperation for the execution of all tasks [31]. Conversely, our method, utilizing a visual planner, empowers the robot to autonomously execute complex daily manipulation tasks, by tracking both 3-DoF trajectories and rough orientations tracking.
@@ III-A Overview
We develop a low-level controller $\pi_{c}$, which allows the robot to track a 3-DoF target trajectory with any forelimb while maintaining stable walking. This is achieved by training a policy whose inputs include the manipulator flag, the current desired end-effector pose, the current target point of the end-effector calculated by the rational Bézier curve, and the next three target points. The output is desired robot joint positions $q^{d}_{m}\in\mathbb{R}^{12}$. This control policy runs at $50$ Hz, and the desired torque $\mathbf{\tau}$ of the robot is obtained through a PD controller from the desired joint positions.
@@ III-C2 Expert Demonstration Collection
As illustrated in Fig. 3, we designed expert trajectory parameters ${}^{o}p$ for different objects and tasks. To enable the robot to learn to manipulate objects placed in different poses, we randomized the positions and yaw axis angles of the objects. The expert trajectory parameters are represented in the object frame, allowing them to accurately follow the object. During the collection of expert demonstrations, the object frame is transformed to the world frame based on the object’s pose.
We gather expert data by having robots follow expertly designed trajectories, utilizing large-scale parallel simulations in IsaacGym [36]. This method enables the rapid collection of a significant volume of data.
The collection time does not increase with the volume of data; it primarily depends on the size of the Video Random Access Memory (VRAM). In this paper, we collect 200 timesteps of expert demonstration data for each task, with 100 trajectories collected per task, taking approximately three minutes.
Despite the collection of expert data in simulations, our method achieves sim-to-real seamless deployment in actual environments, thanks to the point clouds post-process and carefully designed hierarchical structure.
@@ V-A Performance Comparison
Success rate (%) HRL Ours
Press Button $21.67\scriptstyle{\pm 3.51}$ $\mathbf{92.33\scriptstyle{\pm 8.96}}$
Pull Handle $0.00\scriptstyle{\pm 0.00}$ $\mathbf{82.33\scriptstyle{\pm 3.21}}$
Push Door $15.67\scriptstyle{\pm 3.51}$ $\mathbf{85.33\scriptstyle{\pm 5.03}}$
Lift Basket $11.00\scriptstyle{\pm 4.36}$ $\mathbf{59.33\scriptstyle{\pm 12.01}}$
Open Dishwasher $1.33\scriptstyle{\pm 1.15}$ $\mathbf{5.67\scriptstyle{\pm 5.51}}$
Close Dishwasher $4.67\scriptstyle{\pm 1.53}$ $\mathbf{50.33\scriptstyle{\pm 8.14}}$
Pull Objects $8.00\scriptstyle{\pm 2.65}$ $\mathbf{12.67\scriptstyle{\pm 10.79}}$
Twist Valve $10.00\scriptstyle{\pm 4.58}$ $\mathbf{52.33\scriptstyle{\pm 10.21}}$
Shoot Football $3.67\scriptstyle{\pm 2.08}$ $\mathbf{26.00\scriptstyle{\pm 2.00}}$
@@ V-C Sim2real Transfer
We directly implement the trained controller and planner on a real-world Aliengo quadruped robot without the need for further fine-tuning, demonstrating the robustness of our approach. As shown in Fig. 2, the pose of the robot is determined using AprilTag for the coordinate system transformation of trajectory parameters and point clouds. Following this, we applied post-process to the point cloud to ensure alignment with the simulation. In Fig. 1, we illustrate the robot performing tasks like lifting a basket with its forelimbs and pushing the door during locomotion. For demonstrations of all the skills developed through our method, please refer to the website.
@@ VI Conclusion and future works
While we have demonstrated the effectiveness of our method in both simulation and real-world scenarios, there are areas for improvement. (i) The accumulated gap in both two phases results in relatively poor real-world performance. (ii) While our approach to gathering expert demonstrations is significantly efficient, but we need to post-process the point cloud in deployment to align visual observations. (iii) The inference speed limitations of diffusion-based BC hinder task performance, making it challenging for robots to handle dynamic environments.
@@
This study represents a novel effort to master whole-body loco-manipulation, possessing boundless potential. (i) It showcases remarkable scalability, enabling rapid collection of expert data for either scale up or data-mixture with real-world data. (ii) We plan to enhance the planner’s inference speed to equip the framework for more dynamic scenarios.